Find the present value of a perpetuity that produces annual payments of 6500 dollars once per year,starting one year from now.SAssume am effective rate of interest is 10.3 percent.


1
Expert's answer
2021-05-07T10:04:37-0400

According to the questions

Produce annual payments = 6500$ /once per year

 Effective rate of interest = 10.3%


we know that-


"I = \dfrac{P\times R\times T}{100}"


"6500 = \dfrac{P\times 10.3\times 1}{100}"


P = 63,106.7961$